OpenStudy (anonymous):

Estimate the population of bacteria at 9 hours. Time (hours) 0 1 2 3 4 5 6 Population (1000s) 5.1 3.03 1.72 1.17 1.38 2.35 4.08

OpenStudy (anonymous):

@xapproachesinfinity

OpenStudy (xapproachesinfinity):

did you try drawing this!

OpenStudy (xapproachesinfinity):

to see what kind of graph it might be

OpenStudy (anonymous):

No, i don't know where to start. But i do have to show my work

OpenStudy (xapproachesinfinity):

well look for an online plotter put all of those points then post the picture you got here

OpenStudy (anonymous):

ok one second

OpenStudy (xapproachesinfinity):

this looks like a polynomial to me!

OpenStudy (mathmate):

When the numbers are given in equal intervals, another way is to calculate the differences of the given numbers. If the differences are not constant, it means that the function does not represent a straight line. However, you can calculate the differences of the previously calculated differences. If the second difference is constant, the function is a quadratic. You can then continue the sequences to find f(9).

OpenStudy (anonymous):

How do i keep it going to find the answer for 9 hours though

OpenStudy (xapproachesinfinity):

the second difference is constant 0.76 so it really is a quadratic

OpenStudy (anonymous):

so how do we use the quadratic to find the answer? im confused and need to show my work

OpenStudy (xapproachesinfinity):

there is a different method then the sequence f(x)=ax^2+bx+c f(0)=5.1 f(1)=3.03 f(2)=1.17 you got 3 equation in terms of a,b,c solve for them

OpenStudy (xapproachesinfinity):

if you got the equation you plug 9 for f(9)

OpenStudy (anonymous):

A=3.8 B=-2.45 and C=5.1

OpenStudy (xapproachesinfinity):

no no

OpenStudy (anonymous):

I had to find that in the question previous to this one

OpenStudy (anonymous):

Are we doing it differently because we are plugging in the 9?

OpenStudy (xapproachesinfinity):

you have to find a,b , and c plug 0 for one equation 1 for the second 2 doe the third system of three equation

OpenStudy (xapproachesinfinity):

f(0)=5.1=c actually c is good!

OpenStudy (xapproachesinfinity):

how did you find b and a

OpenStudy (xapproachesinfinity):

f(1)=3.03=a+b f(2)=1.72=4a+2b

OpenStudy (xapproachesinfinity):

oh i forgot c a+b+c=3.03 4a+2b+c=1.72

OpenStudy (xapproachesinfinity):

c=5.1

OpenStudy (anonymous):

Ok i've written what we have so far

OpenStudy (xapproachesinfinity):

solve that for a and b

OpenStudy (anonymous):

4a+2b=-3.38?

OpenStudy (xapproachesinfinity):

yes! now you have two equation solve for a and b

OpenStudy (xapproachesinfinity):

i got a=0.38 b=-2.45

OpenStudy (anonymous):

thats what i have

OpenStudy (xapproachesinfinity):

so we have f(x)=0.38x^2-2,45x+5.1

OpenStudy (xapproachesinfinity):

http://prntscr.com/7odnai looks like the shape we have before

OpenStudy (xapproachesinfinity):

plug in 9 for f(9)

OpenStudy (anonymous):

How do i do that though ? thats where i get lost

OpenStudy (xapproachesinfinity):

f(9)=0.38 (9)^2-2.45(9)+5.1

OpenStudy (xapproachesinfinity):

you just replace x by 9

OpenStudy (anonymous):

oh ok, one sec im gonna calculate

OpenStudy (anonymous):

F= 30.78-22.05+5.1

OpenStudy (xapproachesinfinity):

we need final value what is it?

OpenStudy (anonymous):

13.83

OpenStudy (xapproachesinfinity):

yes!

OpenStudy (xapproachesinfinity):

that's the number of bacteria after 9hours in thousands

OpenStudy (anonymous):

So thats our final answer 13.83? :)